OpenStudy (anonymous):

Change the given equation into slope-intercept form. 3x – 4y = 12

OpenStudy (johnweldon1993):

Slope intercept form is y=mx+b so 3x - 4y = 12 -4y = 12 - 3x y = -3 + 3/4x y = 3/4x - 3

OpenStudy (anonymous):

thanks!

OpenStudy (johnweldon1993):

Anytime :)